# literature.bib

### v4

@@ 5. Application to fluid dynamics
In this subsection, we apply our results to an abstract fluid dynamics model considered in several earlier works. We closely follow the presentation of [AVsurvey, chueshovmillet10] and focus on what the large deviation principle of Theorem 2.6 becomes in this setting. Afterwards, we specialize to the Navier–Stokes equations with gradient noise to make our results even more concrete.
@@ 5.1. Abstract model
[…]
@@
In [AVsurvey, Th. 7.10] it is shown that Assumption 5.1 is satisfied, noting that the arguments also work for the time-dependent setting. Thus well-posedness follows from Theorem 2.3 and the large deviation principle follows from Theorem 2.6. ∎
@@ 5.2. LDP for Navier–Stokes equations with gradient noise
For simplicity we only consider the case of Itô noise. For details on Stratonovich noise, see [AV24SNS, App. A]. We follow the presentation of [AVsurvey, §7.3.4].
@@
(5.3) $\,\mathrm{d}Y^{\varepsilon}&=\big{[}\nu\Delta Y^{\varepsilon}-(Y^{\varepsilon}\cdot\nabla)Y^{\varepsilon}-\nabla P^{\varepsilon}\big{]}\,\,\mathrm{d}t\\ &\qquad+\sqrt{\varepsilon}\sum_{n\geq 1}\big{[}(b_{n}\cdot\nabla)Y^{\varepsilon}+g_{n}(\cdot,Y^{\varepsilon})-\nabla\widetilde{P}_{n}^{\varepsilon}\big{]}\,\,\mathrm{d}W_{t}^{n},\\ {\text{div}}\,Y^{\varepsilon}&=0,\\ Y^{\varepsilon}&=0\ \text{on $\partial\mathcal{O}$},\\ Y^{\varepsilon}(0,\cdot)&=u_{0}.\end{aligned}\right.$
@@
As in [AVsurvey, §7.3.4], we can use the Helmholtz projection $\mathbb{P}$ to rewrite (5.3) as (5.1). To this end, let $\mathcal{U}=\ell^{2}$ with standard basis $(e_{n})_{n\geq 1}$ and let
$H=\mathbb{L}^{2}(\mathcal{O}),\ \ V=\mathbb{H}^{1}_{0}(\mathcal{O})=H^{1}_{0}(\mathcal{O};{\mathbb{R}}^{2})\cap\mathbb{L}^{2}(\mathcal{O})\ \ \text{and}\ \ V^{*}\coloneqq\mathbb{H}^{-1}(\mathcal{O})=(\mathbb{H}^{1}_{0}(\mathcal{O}))^{*},$
@@
$A_{0}=-\nu\mathbb{P}\Delta,\ \ \Phi(u,v)=-\mathbb{P}{\text{div}}[u\otimes v],\ (B_{0}u)e_{n}=\mathbb{P}[(b_{n}\cdot\nabla)u],\ \text{and}\ G(u)e_{n}=\mathbb{P}g_{n}(\cdot,u).$
@@
(5.4) $\,\mathrm{d}u^{\psi}&=\big{[}\nu\Delta u^{\psi}-\mathbb{P}{\text{div}}(u^{\psi}\otimes u^{\psi})\big{]}\,\,\mathrm{d}t+\sqrt{\varepsilon}\sum_{n\geq 1}\big{(}\mathbb{P}[(b_{n}\cdot\nabla)u^{\psi}]+\mathbb{P}g_{n}(\cdot,u^{\psi})\big{)}\psi_{n},\\ u^{\psi}&=0\ \text{on $\partial\mathcal{O}$},\\ u^{\psi}(0,\cdot)&=u_{0}.\end{aligned}\right.$
In [AVsurvey, §7.3.4] it is verified that Assumption 5.1 is fulfilled for the above setting. Thus we obtain the next result immediately from Theorem 5.2.
@@
The following special case of a chain rule from [pardoux] is useful, since it applies to $L^{2}(0,T;V^{*})+L^{1}(0,T;H)$-valued integrands.
@@
[pardoux, Lem. 2.2 p. 30] Let $(V,H,V^{*})$ be a Gelfand triple of Hilbert spaces. Let $x\in H$, $u\in C([0,T];H)\cap L^{2}(0,T;V)$ and $v\in L^{2}(0,T;V^{*})+L^{1}(0,T;H)$ be such that
@@
Note that $u\in L^{2}(0,T;V)\cap L^{\infty}(0,T;H)$ and $v\in L^{1}(0,T;V^{*})$. Thus by (A.1), $u$ is weakly differentiable with $u^{\prime}=v$ a.e. on $[0,T]$, see [HNVWvolume1, Lem. 2.5.8]. Hence, $u^{\prime}=v\in L^{2}(0,T;V^{*})+L^{1}(0,T;H)$. Also, (A.1) implies absolute continuity of $u:[0,T]\to V^{*}$. Now [pardoux, Lem. 2.2 p. 30, $p=2$] gives $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|u(t)\|_{H}^{2}=2\langle u^{\prime}(t),u(t)\rangle=2\langle v(t),u(t)\rangle$ a.e., proving (A.2). ∎
Stochastic versions of the chain rule, or Itô formula, are also given in [pardoux]. The following special case is suited for random, $L^{2}(0,T;V^{*})+L^{1}(0,T;H)$-valued integrands. We recall that the class of integrable processes for a $U$-cylindrical Brownian motion (Definition 4.1) is given by
(A.3) $\mathcal{N}(0,T)\coloneqq$$\Bigl{\{}\Phi\colon[0,T]\times\Omega\to\mathcal{L}_{2}(U,H):\Phi\text{ strongly progressively measurable, }$,
@@
[pardoux, Th. 3.1 p. 57, Th. 3.3 p. 59] Let $(V,H,V^{*})$ be a Gelfand triple of Hilbert spaces and let $(\Omega,\mathcal{F},\mathbb{P},(\mathcal{F}_{t})_{t\in{\mathbb{R}}_{+}})$ be a filtered probability space. Suppose that
@@
    $v\in L^{0}(\Omega;L^{1}(0,T;H))+L^{0}(\Omega;L^{2}(0,T;V^{*}))$, $v$ is adapted,
@@
(A.4) $\|u(t)\|_{H}^{2}=\|u_{0}\|_{H}^{2}+2\int_{0}^{t}\langle v(s),u(s)\rangle\,\mathrm{d}s+2\int_{0}^{t}\langle u(s),\Phi(s)\,\mathrm{d}W(s)\rangle+\int_{0}^{t}|\!|\!|\Phi(s)|\!|\!|_{H}^{2}\,\mathrm{d}s.$
Finally, we relate the $U$-cylindrical Brownian motion $W$ of Definition 4.1 to the ${\mathbb{R}}^{\infty}$-Brownian motion $\tilde{W}$ of Definition 4.2, as well as their stochastic integrals constructed in [NVW15] and [liurockner15], respectively.
An ${\mathbb{R}}^{\infty}$-Brownian motion $\tilde{W}=((\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$ in $U$ corresponds to a Wiener process $\tilde{W}_{1}$ in a larger space $U_{1}$, with trace class covariance. That is, for any Hilbert-Schmidt embedding $J\colon U\hookrightarrow U_{1}$, the $U_{1}$-valued process given by
@@
defines a $Q_{1}$-Wiener process on $U_{1}$, with $Q_{1}\coloneqq JJ^{*}\in\mathcal{L}(U_{1},U_{1})$ nonnegative definite, symmetric and of trace class [liurockner15, Prop. 2.5.2].
It is well-known that $\mathcal{N}(0,T)$ from (A.3) is the class of integrable processes for both $\tilde{W}$ and $W$, see [liurockner15, p. 52, p. 53], [AV22nonlinear1, Prop. 2.13] and the proof in [NVW15, p. 306, §5.4 $(p=0)$]. The next proposition relates the stochastic integrals corresponding to $\tilde{W}$ and $W$.
@@
For any $U$-cylindrical Brownian motion $W\in\mathcal{L}(L^{2}({\mathbb{R}}_{+};U),L^{2}(\Omega))$ and any orthonormal basis $(e_{k})_{k\in{\mathbb{N}}}$ of $U$, there exists an ${\mathbb{R}}^{\infty}$-Brownian motion $\tilde{W}=((\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$ with
(A.6) $W(\mathbbm{1}_{(0,t]}\otimes e_{k})=\beta_{k}(t)\text{\quad in }L^{2}(\Omega),\quad\text{ for all }k\in{\mathbb{N}}\text{ and }t\in{\mathbb{R}}_{+}.$
@@
Reversely, given an ${\mathbb{R}}^{\infty}$-Brownian motion $\tilde{W}=((\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$, there exists a unique $U$-cylindrical Brownian motion $W\in\mathcal{L}(L^{2}({\mathbb{R}}_{+};U),L^{2}(\Omega))$ that satisfies (A.6).
@@
with $\tilde{W}_{1}$ as in (A.5). Here, the integral on the left-hand side is the one constructed in [NVW15] and the middle and right integral are those constructed in [liurockner15, §2.3, §2.5].
[…]

### v5

@@ 5. Application to fluid dynamics
In this subsection, we apply our results to an abstract fluid dynamics model considered in several earlier works. We closely follow the presentation of [undefb, undefk] and focus on what the large deviation principle of Theorem 2.6 becomes in this setting. Afterwards, we specialize to the Navier–Stokes equations with gradient noise to make our results even more concrete.
@@ 5.1. Abstract model
(5.1) $\left\{\begin{aligned} &\,\mathrm{d}Y^{\varepsilon}(t)+A_{0}(t)Y^{\varepsilon}\,\,\mathrm{d}t=\Phi(Y^{\varepsilon}(t),Y^{\varepsilon}(t))\,\,\mathrm{d}t+\sqrt{\varepsilon}\big(B_{0}(t)Y^{\varepsilon}(t)+G(t,Y^{\varepsilon}(t))\big)\,\,\mathrm{d}W(t),\\ &u(0)=x.\end{aligned}\right$.
@@
In [undefb, Th. 7.10] it is shown that Assumption 5.1 is satisfied, noting that the arguments also work for the time-dependent setting. Thus well-posedness follows from Theorem 2.3 and the large deviation principle follows from Theorem 2.6. ∎
@@ 5.2. LDP for Navier–Stokes equations with gradient noise
For simplicity we only consider the case of Itô noise. For details on Stratonovich noise, see [undefd, App. A]. We follow the presentation of [undefb, §7.3.4].
@@
(5.3) $\left\{\begin{aligned} &\,\mathrm{d}Y^{\varepsilon}=\big[\nu\Delta Y^{\varepsilon}-(Y^{\varepsilon}\cdot\nabla)Y^{\varepsilon}-\nabla P^{\varepsilon}\big]\,\,\mathrm{d}t+\sqrt{\varepsilon}\sum_{n\geq 1}\big[(b_{n}\cdot\nabla)Y^{\varepsilon}+g_{n}(\cdot,Y^{\varepsilon})-\nabla\widetilde{P}_{n}^{\varepsilon}\big]\,\,\mathrm{d}W_{t}^{n},\\ &{\text{div}}\,Y^{\varepsilon}=0,\\ &Y^{\varepsilon}=0\quad\text{on $\partial\mathcal{O}$},\\ &Y^{\varepsilon}(0,\cdot)=u_{0}.\end{aligned}\right$.
@@
As in [undefb, §7.3.4], we can use the Helmholtz projection $\mathbb{P}$ to rewrite (5.3) as (5.1). To this end, let $\mathcal{U}=\ell^{2}$ with standard basis $(e_{n})_{n\geq 1}$ and let
$H=\mathbb{L}^{2}(\mathcal{O}),\quad V=\mathbb{H}^{1}_{0}(\mathcal{O})=H^{1}_{0}(\mathcal{O};{\mathbb{R}}^{2})\cap\mathbb{L}^{2}(\mathcal{O}),\quad V^{*}\coloneqq\mathbb{H}^{-1}(\mathcal{O})=(\mathbb{H}^{1}_{0}(\mathcal{O}))^{*},$
@@
$A_{0}=-\nu\mathbb{P}\Delta,\quad\Phi(u,v)=-\mathbb{P}{\text{div}}[u\otimes v],\quad(B_{0}u)e_{n}=\mathbb{P}[(b_{n}\cdot\nabla)u],\quad G(u)e_{n}=\mathbb{P}g_{n}(\cdot,u).$
@@
(5.4) $\left\{\begin{aligned} &\,\mathrm{d}u^{\psi}=\big[\nu\Delta u^{\psi}-\mathbb{P}{\text{div}}(u^{\psi}\otimes u^{\psi})\big]\,\,\mathrm{d}t+\sum_{n\geq 1}\big(\mathbb{P}[(b_{n}\cdot\nabla)u^{\psi}]+\mathbb{P}g_{n}(\cdot,u^{\psi})\big)\psi_{n},\\ &u^{\psi}=0\quad\text{on $\partial\mathcal{O}$},\\ &u^{\psi}(0,\cdot)=u_{0}.\end{aligned}\right$.
In [undefb, §7.3.4] it is verified that Assumption 5.1 is fulfilled for the above setting. Thus we obtain the next result immediately from Theorem 5.2.
@@
The following special case of a chain rule from [undefag] is useful, since it applies to $L^{2}(0,T;V^{*})+L^{1}(0,T;H)$-valued integrands.
@@
[undefag, Lem. 2.2 p. 30] Let $(V,H,V^{*})$ be a Gelfand triple of Hilbert spaces. Let $x\in H$, $u\in C([0,T];H)\cap L^{2}(0,T;V)$ and $v\in L^{2}(0,T;V^{*})+L^{1}(0,T;H)$ be such that
@@
Note that $u\in L^{2}(0,T;V)\cap L^{\infty}(0,T;H)$ and $v\in L^{1}(0,T;V^{*})$. Thus by (A.1), $u$ is weakly differentiable with $u^{\prime}=v$ a.e. on $[0,T]$, see [undeft, Lem. 2.5.8]. Hence, $u^{\prime}=v\in L^{2}(0,T;V^{*})+L^{1}(0,T;H)$. Also, (A.1) implies absolute continuity of $u:[0,T]\to V^{*}$. Now [undefag, Lem. 2.2 p. 30, $p=2$] gives $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|u(t)\|_{H}^{2}=2\langle u^{\prime}(t),u(t)\rangle=2\langle v(t),u(t)\rangle$ a.e., proving (A.2). ∎
Stochastic versions of the chain rule, or Itô formula, are also given in [undefag]. The following special case is suited for random, $L^{2}(0,T;V^{*})+L^{1}(0,T;H)$-valued integrands. We recall that the class of integrable processes for a $U$-cylindrical Brownian motion (Definition 4.1) is given by
(A.3) $\mathcal{N}(0,T)\coloneqq$ $\Bigl\{\Phi\colon[0,T]\times\Omega\to\mathcal{L}_{2}(U,H):\Phi\text{ strongly progressively measurable, }$
@@
[undefag, Th. 3.1 p. 57, Th. 3.3 p. 59] Let $(V,H,V^{*})$ be a Gelfand triple of Hilbert spaces and let $(\Omega,\mathcal{F},\mathbb{P},(\mathcal{F}_{t})_{t\in{\mathbb{R}}_{+}})$ be a filtered probability space. Suppose that
@@
    $\in L^{0}(\Omega;L^{1}(0,T;H))+L^{0}(\Omega;L^{2}(0,T;V^{*})), vv$ is adapted,
@@
[…]
@@
defines a $Q_{1}$-Wiener process on $U_{1}$, with $Q_{1}\coloneqq JJ^{*}\in\mathcal{L}(U_{1},U_{1})$ nonnegative definite, symmetric and of trace class [undefab, Prop. 2.5.2].
It is well-known that $\mathcal{N}(0,T)$ from (A.3) is the class of integrable processes for both $\tilde{W}$ and $W$, see [undefab, p. 52, p. 53], [undef, Prop. 2.13] and the proof in [undefad, p. 306, §5.4 $(p=0)$]. The next proposition relates the stochastic integrals corresponding to $\tilde{W}$ and $W$.
@@
[…]
@@
[…]
@@
with $\tilde{W}_{1}$ as in (A.5). Here, the integral on the left-hand side is the one constructed in [undefad] and the middle and right integral are those constructed in [undefab, §2.3, §2.5].